# Don’t Make It Up: Preserving Ignorance Awareness in LLM Fine-Tuning

### v2

@@ 7 Related Works
Continual learning for LLMs has emerged as a critical area of research, motivated by the need to efficiently incorporate new knowledge without catastrophic forgetting of prior knowledge. Traditional approaches, such as rehearsal-based methods (lopez2017gradient, ) and parameter isolation techniques (serra2018overcoming, ), have been adapted to the LLM setting, but face unique challenges due to issues surrounding their scalability and sensitivity. Recent work has explored modular architectures and adapter-based methods to localize task-specific updates and reduce interference with general knowledge (wang2024localizing, ). Others have proposed continual learning through task arithmetic. (ilharco2022editing, ) pioneered the approach of training each task separately using LoRA and subsequently merging tasks via task arithmetic, as opposed to sequential task training. Despite these advances, preserving key alignment behaviors, such as factual accuracy and refusal to answer unverifiable or harmful prompts, remains difficult in a continual learning setting. Our work builds on this line of research by introducing a continual fine-tuning strategy that preserves safety-aligned behaviors while maintaining adaptability to novel data distributions, addressing the critical problem of alignment retention in LLMs.
@@ 8 Conclusion
We formalize the notion of ‘ignorance awareness’ in LLMs and propose SEAT, a simple and principled method for robust LLM fine-tuning that excels at incorporating new knowledge while preserving the model’s pre-aligned ability to faithfully express ignorance towards unseen data. Through comprehensive empirical analysis, we demonstrate SEAT ’s effectiveness across various training configurations, as well as the complementary and essential roles of its two components in maintaining model’s calibrated response behavior.
@@ Appendix A Proofs
A.1 Proof of Proposition 1
Proposition 1
@@ A.1 Proof of Proposition 1
A decoder-only transformer model is a finite composition of primitives. Using Llama3 [3] as a proxy, we list its modules, the formula implemented and its smoothness class below.
@@ A.2 Proof of Proposition 2
$L_{\ell}(K)\;:=\;\sup_{\theta\in K}\;\bigl{\|}\nabla_{\theta}R_{\ell}(\theta;x)\bigr{\|}_{\operatorname{op}}\;<\;\infty.$
@@
A.3 Proof of Theorem 1
Let $\mathcal{U}\subseteq\{1,\dots,P\}$ be the trainable coordinates and $\mathcal{F}=\mathcal{U}^{\mathrm{c}}$ be the frozen ones. Define sparse fine-tuning as $\theta^{\prime}\;=\;\theta\;-\;\eta\,M\,\nabla_{\theta}L(\theta)$, where $M$ is the mask matrix.
@@ A.3 Proof of Theorem 1
Theorem 1
@@
$\bigl{\lVert}M\nabla_{\theta}\mathcal{L}(\theta)\bigr{\rVert}\leq\bigl{\lVert}\nabla_{\theta}\mathcal{L}(\theta)\bigr{\rVert}$
with equality if and only if the gradient has no component in any frozen coordinate: $\nabla_{\theta}L(\theta)]_{i}=0$ for all $i\in\mathcal{F}$.
@@
Now, we show the basic primitives used in transformers are both input and parameter-Lipschitz bounded. Throughout let $\|\cdot\|$ be the Euclidean norm and $\|\cdot\|_{\operatorname{op}}$ the corresponding operator norm.
Lemma 3 (Input Lipschitz constants).
For the basic primitives used in transformers, the following bounds hold for every $x\in\mathbb{R}^{d}$:
$\|x\mapsto Wx\|_{\operatorname{op}}$ $=\|W\|_{\operatorname{op}},$
$\|x\mapsto\operatorname{RoPE}(x)\|_{\operatorname{op}}$ $=1,$
$\|x\mapsto\sigma(x)\|_{\operatorname{op}}$ $\leq 1,$
$\|\nabla_{x}\operatorname{SwiGLU}(x)\|_{\operatorname{op}}$ $\leq 2\|x\|_{\infty},$
$\bigl{\|}x\mapsto\operatorname{RMSNorm}_{\gamma,\varepsilon}(x)\bigr{\|}_{\operatorname{op}}$ $\leq\|\gamma\|_{\infty},$
$\text{and}\quad\|x\mapsto x+F(x)\|_{\operatorname{op}}$ $\leq 1+\|F\|_{\operatorname{op}}\quad\text{for any map }F.$
@@
1. Linear map
The Jacobian equals $W$; its spectral norm is $\|W\|_{\operatorname{op}}$.
2. RoPE
Rotary position encoding multiplies each 2-slice $(x_{2k},x_{2k+1})$ by an orthogonal $2\times 2$ rotation matrix. The full Jacobian is block-diagonal with orthogonal blocks, hence has spectral norm 1.
3. Soft-max
At $z\in\mathbb{R}^{d}$, the Jacobian is
$J_{ij}(z)=\sigma_{i}(z)\bigl{(}\delta_{ij}-\sigma_{j}(z)\bigr{)}.$
This symmetric doubly-stochastic matrix has eigenvalues in $[0,1]$; therefore $\|J(z)\|_{\operatorname{op}}\leq 1$ for every $z$.
4. SwiGLU
Write the input as $x=(u,v)\in\mathbb{R}^{2d}$. Component-wise, $f_{i}(u,v)=\operatorname{Swish}(u_{i})\,v_{i}$ with $\operatorname{Swish}(t)=t\sigma(t)$. Since
$\operatorname{Swish}^{\prime}(t)=\sigma(t)+t\sigma(t)\bigl{(}1-\sigma(t)\bigr{)}$
attains its global maximum $\beta\approx 1.09984<1.1$,
$|\partial_{u_{i}}f_{i}|\leq\beta|v_{i}|,\qquad|\partial_{v_{i}}f_{i}|\leq|u_{i}|.$
Each $2\times 1$ row of the Jacobian is therefore bounded by $\sqrt{\beta^{2}+1}\,\|x\|_{\infty}<2\,\|x\|_{\infty}$. The rows are orthogonal, so the full spectral norm obeys the same bound.
5. RMSNorm
Let $g(x)=\|x\|^{2}/d+\varepsilon$. Then
$\nabla_{x}\operatorname{RMSNorm}_{\gamma,\varepsilon}(x)=\gamma\Bigl{(}g(x)^{-1/2}I_{d}-\tfrac{1}{2d}g(x)^{-3/2}xx^{\top}\Bigr{)}.$
The first term has norm $\|\gamma\|_{\infty}g(x)^{-1/2}\leq\|\gamma\|_{\infty}$. The rank-1 correction has smaller norm, so the whole Jacobian is bounded by $\|\gamma\|_{\infty}$.
6. Residual connection
For any $x,y\in\mathbb{R}^{d}$,
$\|x+F(x)-y-F(y)\|$ $\leq\|x-y\|+\|F(x)-F(y)\|$
$\leq(1+\|F\|_{\operatorname{op}})\|x-y\|.$
Lemma 4 (Parameter Lipschitz constants).
For the basic primitives used in transformers, there exists a constant $c_{\mathrm{prim}}>0$ (depending only on architecture hyperparameters and the fixed offset $\varepsilon>0$) such that
$\|\nabla_{\theta}f_{\theta}(x)\|_{\operatorname{op}}\leq c_{\mathrm{prim}}\,\bigl{(}1+\|x\|\bigr{)}$
for every admissible $(\theta,x)\in\Theta\times\mathbb{R}^{d}$. Consequently every primitive map $\theta\mapsto f_{\theta}(x)$ is Lipschitz with constant growing at most linearly in $\|x\|$.
Proof.
1. Linear map
Let $\theta=\operatorname{vec}W\in\mathbb{R}^{d\times m}$, a first-order variation $\delta\theta=\operatorname{vec}\bigl{(}\delta W\bigr{)}$ produces $\delta f=\delta W\,x$. Hence $\nabla_{\theta}f_{\theta}(x)=x^{\top}\otimes I_{d}\in\mathbb{R}^{d\times(d\times m)}.$
Since $\|A\otimes B\|_{\operatorname{op}}=\|A\|_{\operatorname{op}}\|B\|_{\operatorname{op}}$, $\|x^{\top}\|_{\operatorname{op}}=\|x\|$ and $\|I_{d}\|_{\operatorname{op}}=1$, we show $\bigl{\|}\nabla_{\theta}f_{\theta}(x)\bigr{\|}_{\operatorname{op}}=\|x\|\leq 1+\|x\|.$ and, thus, $c_{\mathrm{lin}}:=1$.
2. RoPE
RoPE is parameter-free. Hence $\nabla_{\theta}f_{\theta}(x)\equiv 0$ and $c_{\mathrm{RoPE}}:=0$.
3. Soft-max
The canonical implementation of soft-max has no learnable parameters, so again $\nabla_{\theta}f_{\theta}(x)\equiv 0$ and $c_{\sigma}:=0$.
4. SwiGLU
Let $\theta=(\operatorname{vec}W_{1},b_{1},\operatorname{vec}W_{2},b_{2})\in\mathbb{R}^{d_{1}d+d_{1}+dd_{1}+d},$ where $W_{1}\in\mathbb{R}^{d_{1}\times d},\;W_{2}\in\mathbb{R}^{d\times d_{1}}$.
Derivatives w.r.t. $(W_{2},b_{2})$
$\partial_{W_{2}}f_{\theta}(x)$ $=\operatorname{SwiGLU}(W_{1}x+b_{1})$
$\Longrightarrow\quad\|\partial_{W_{2}}f_{\theta}(x)\|_{\operatorname{op}}$ $\leq\|W_{1}x+b_{1}\|,$
$\partial_{b_{2}}f_{\theta}(x)$ $=I_{d}$
$\Longrightarrow\quad\|\partial_{b_{2}}f_{\theta}(x)\|_{\operatorname{op}}$ $=1.$
Because $\|W_{1}x+b_{1}\|\leq\|W_{1}\|_{\operatorname{op}}\|x\|+\|b_{1}\|$, there exists a constant $c_{1}$ (the maximum of $\|W_{1}\|_{\operatorname{op}}$ and $\|b_{1}\|$) such that
$\|(\partial_{W_{2}}f,\partial_{b_{2}}f)\|_{\operatorname{op}}\leq c_{1}(1+\|x\|).$
Derivatives w.r.t. $(W_{1},b_{1})$
Let $a=W_{1}x+b_{1}\in\mathbb{R}^{2d_{1}}$ (split into gates $u,v\in\mathbb{R}^{d_{1}}$). Lemma 3 gives
$\|\nabla_{a}\operatorname{SwiGLU}(a)\|_{\operatorname{op}}\leq 2\|a\|_{\infty}.$
Hence
$\partial_{W_{1}}f_{\theta}(x)$ $=W_{2}\,\nabla_{a}\operatorname{SwiGLU}(a)\,x^{\top}$
$\partial_{b_{1}}f_{\theta}(x)$ $=W_{2}\,\nabla_{a}\operatorname{SwiGLU}(a).$
Bounding $\|a\|_{\infty}$:
$\|a\|_{\infty}\leq\|W_{1}\|_{\operatorname{op}}\|x\|+\|b_{1}\|_{\infty}.$
Taking operator norms,
$\|\partial_{W_{1}}f_{\theta}(x)\|_{\operatorname{op}}$ $\leq\|W_{2}\|_{\operatorname{op}}\cdot 2\|a\|_{\infty}\cdot\|x\|$
$\leq 2\|W_{2}\|_{\operatorname{op}}\bigl{(}\|W_{1}\|_{\operatorname{op}}\|x\|+\|b_{1}\|_{\infty}\bigr{)}\|x\|,$
$\|\partial_{b_{1}}f_{\theta}(x)\|_{\operatorname{op}}$ $\leq 2\|W_{2}\|_{\operatorname{op}}\|a\|_{\infty}.$
Both are bounded by $c_{2}(1+\|x\|)$ with
$c_{2}=2\|W_{2}\|_{\operatorname{op}}\max\{\|W_{1}\|_{\operatorname{op}},\|b_{1}\|_{\infty},1\}.$
Thus, the combined $c_{\mathrm{Swi}}:=\max(c_{1},c_{2})$.
5. RMSNorm
Let $\theta=(\gamma,\beta)\in\mathbb{R}^{2d}$ and $g(x)=\|x\|^{2}/d+\varepsilon$.
$\partial_{\gamma}f_{\theta}(x)$ $=\operatorname{diag}\left(\frac{x}{\sqrt{g(x)}}\right)$
$\partial_{\beta}f_{\theta}(x)$ $=I_{d}$
$\Longrightarrow\quad\bigl{\|}\partial_{\gamma}f_{\theta}(x)\bigr{\|}_{\operatorname{op}}$ $\leq\frac{\|x\|}{\sqrt{d\varepsilon}}$
$\bigl{\|}\partial_{\beta}f_{\theta}(x)\bigr{\|}_{\operatorname{op}}$ $=1.$
Thus, $c_{\mathrm{RMS}}:=\max\left(1,\frac{1}{\sqrt{d\varepsilon}}\right).$
A.4 Proof of Theorem 2
Theorem 2
Assume the training trajectory $\{\theta_{t}\}_{t\geq 0}\subset\Theta$ remains in a compact set $K$. Let $\ell$ be any layer and put $L_{\ell}:=L_{\ell}(K)$ from Proposition 2. For a deterministic gradient step $\theta^{\prime}=\theta-\eta\nabla_{\theta}\mathcal{L}(\theta)$ with step-size $\eta>0$ we have
$\bigl{\|}R_{\ell}(\theta^{\prime})-R_{\ell}(\theta)\bigr{\|}\;\leq\;\eta\,L_{\ell}\,\bigl{\|}\nabla_{\theta}\mathcal{L}(\theta)\bigr{\|}.$
Proof.
Let $\gamma(t)=\theta+t(\theta^{\prime}-\theta)$ for $t\in[0,1]$. By the fundamental theorem of calculus for curves in $\mathbb{R}^{m}$
@@ A.4 Proof of Theorem 2
The segment $\gamma([0,1])\subset K$ by assumption, hence the supremum is $\leq L_{\ell}$. Finally $\|\theta^{\prime}-\theta\|=\eta\|\nabla_{\theta}\mathcal{L}(\theta)\|$, yielding the deterministic bound. ∎
Corollary 1
Assume the mask $M$ is drawn independently of the gradient, freezing each coordinate with probability $\alpha\in[0,1)$. For any $g\in\mathbb{R}^{P}$,
$\mathbb{E}\bigl{[}\|Mg\|^{2}\bigr{]}$ $=(1-\alpha)\,\|g\|^{2},\quad\text{and}$
$\mathbb{E}\bigl{[}\|Mg\|\bigr{]}$ $\leq\sqrt{1-\alpha}\,\|g\|.$
Proof.
Since $M$ is diagonal, $\|Mg\|^{2}=\sum_{i}m_{i}g_{i}^{2}$ and $\mathbb{E}m_{i}=1-\alpha$, giving the first identity. The second line follows from Jensen’s inequality $\mathbb{E}\|Mg\|\leq\sqrt{\mathbb{E}\|Mg\|^{2}}$. ∎
Corollary 2
If $\mathcal{U}_{1}\subseteq\mathcal{U}_{2}$, then for every $g\in\mathbb{R}^{P}$,
$\|{M_{\mathcal{U}_{1}}g\|}\;\leq\;\|{M_{\mathcal{U}_{2}}g}\|\;\leq\;\|{g}\|.$
Proof.
Because $M_{\mathcal{U}_{1}}=M_{\mathcal{U}_{1}}M_{\mathcal{U}_{2}}$ and both masks are orthogonal projections, Lemma 2 gives $\|{M_{\mathcal{U}_{1}}g}\|\leq\|{M_{\mathcal{U}_{2}}g}\|\leq\|{g}\|$. ∎
Corollary 3 (Stochastic gradient step).
If instead a stochastic gradient $g(\theta,\xi)$ is used, then taking expectations (over $\xi$) gives
$\mathbb{E}\bigl{[}\|R_{\ell}(\theta^{\prime})-R_{\ell}(\theta)\|\bigr{]}\leq\eta\,L_{\ell}\,\mathbb{E}\bigl{[}\|g(\theta,\xi)\|\bigr{]}.$
Proof.
The stochastic inequality follows by taking expectations and Jensen’s inequality. ∎
Corollary 4 (Adam-type steps).
Suppose the preconditioner $\hat{v}_{t}^{-1/2}$ in an Adam-type update $\theta^{\prime}=\theta-\eta_{t}\,\hat{v}_{t}^{-1/2}\!\odot m_{t}$ is almost surely bounded by a constant $c>0$ (coordinate-wise). Then
$\mathbb{E}\bigl{[}\|R_{\ell}(\theta^{\prime})-R_{\ell}(\theta)\|\bigr{]}\;\leq\;\eta_{t}\,c\,L_{\ell}\,\mathbb{E}\bigl{[}\|m_{t}\|\bigr{]}.$
Proof.
Replace $\theta^{\prime}-\theta$ in the previous proof by $\eta_{t}\,\hat{v}_{t}^{-1/2}\!\odot m_{t}$ and use $\|\hat{v}_{t}^{-1/2}\!\odot m_{t}\|\leq c\,\|m_{t}\|$. ∎
Remarks.
If weight-decay and gradient-clipping are in force, they empirically keep the trajectory in a bounded ball; mathematically this is captured by the compact-set hypothesis in Proposition 2. Lemma 3 is useful for bounding $\|R_{\ell}(\theta;x)\|$ with respect to $x$, whereas Lemma 4 underlies explicit numerical estimates of $L_{\ell}$.
A.5 Proof of Theorem 3
Lemma 5 (Scoring function Lipschitz constants).
Let $S:\{0,1\}\times(0,1)\to\mathbb{R}$ be the binary cross-entropy loss defined by $S(b,p):=-b\log p-(1-b)\log(1-p)$, for binary state of known or unknown by the LLM $b\in\{0,1\}$ and predicted probabilities $p\in(0,1)$. Then for any fixed $\delta\in(0,\tfrac{1}{2})$, the function $S$ satisfies the uniform Lipschitz property:
$\bigl{|}S(b,p)-S(b,p^{\prime})\bigr{|}\leq L_{\delta}\cdot|p-p^{\prime}|,$
$\forall b\in\{0,1\},\;p,p^{\prime}\in[\delta,1-\delta],$
where the Lipschitz constant is $L_{\delta}:=\max\left\{\frac{1}{\delta},\frac{1}{1-\delta}\right\}$.
Proof.
When $b=1$,
$|S^{\prime}(p)|=\frac{1}{p}\leq\frac{1}{\delta},\quad\forall p\in[\delta,1-\delta].$
Similarly, when $b=0$,
$|S^{\prime}(p)|=\frac{1}{1-p}\leq\frac{1}{1-\delta},\quad\forall p\in[\delta,1-\delta].$
Combining both cases, we have:
$\sup_{b\in\{0,1\},\;p\in[\delta,1-\delta]}\left|\frac{d}{dp}f(b,p)\right|\;\leq\;\max\left\{\frac{1}{\delta},\frac{1}{1-\delta}\right\}=L_{\delta}.$
Applying the Mean Value Theorem, we establish that $S$ is Lipschitz continuous with constant $L_{\delta}$ over the interval $\delta,1-\delta]$.
Theorem 3
For a proper Bernoulli scoring rule $S$ that fulfills the uniform $L_{\delta}$-Lipschitz property and assume the ignorance score functional $f_{\theta}:\mathbb{R}^{d}\!\to\![0,1]$ is $C_{f}$-Lipschitz bound, the change of ignorance awareness satisfies the bound
$\bigl{\lVert}\Delta_{\text{\emph{AoI}},S}(\theta\!\to\!\theta^{\prime})\bigl{\lVert}\;\leq\;L_{\delta}\,C_{f}\,\varepsilon$
Proof.
We begin by expanding the definition of the change of ignorance awareness:
$\Delta_{\mathrm{IA}}(\theta\to\theta^{\prime})$ $=\mathbb{E}\left[S(I,f(\theta^{\prime};Q))-S(I,f(\theta;Q)))\right].$
Apply the triangle inequality to the absolute value, we get:
$\bigl{\lVert}\Delta_{\mathrm{IA}}(\theta\to\theta^{\prime})\bigl{\lVert}$ $\leq\mathbb{E}\left[\bigl{\lVert}S(I,f(\theta^{\prime};Q))-S(I,f(\theta;Q))\bigl{\lVert}\right].$
Now, apply Lipschitz continuity of the scoring rule $S$ (refer to Lemma 5) in its second argument:
$\bigl{\lVert}S(I,f(\theta^{\prime};Q))-S(I,f(\theta;Q))\bigl{\lVert}\leq L_{\delta}\cdot\bigl{\lVert}f(\theta^{\prime};Q)-f(\theta;Q)\bigl{\lVert}$
Assume the Lipschitz continuity of the score functional $f$ with constant $C_{f}$ (and rewrite its argument as $R(\theta)$ represents the residual stream activation of a model parameterized by $\theta$ in response to query $Q$), we obtain:
$\bigl{\lVert}f(R(\theta^{\prime};Q))-f(R(\theta;Q))\bigl{\lVert}\leq C_{f}\cdot\bigl{\lVert}R(\theta^{\prime};Q)-R(\theta;Q)\bigl{\lVert}.$
Note that this assumption is justified by the observation that a well-aligned language model should exhibit stable estimates of ignorance awareness under small perturbations of its internal representations. Empirical studies support this assumption, showing that activation regions associated with ignorance states tend to be substantially broader than those corresponding to finely localized, precise knowledge [22].
Combining the above, we obtain:
$\bigl{\lVert}S(I,f(\theta^{\prime};Q))-S(I,f(\theta;Q))\bigl{\lVert}\leq L_{\delta}\cdot C_{f}\cdot\varepsilon,$
where $\varepsilon$ is the residual stream activation displacement $\bigl{\lVert}R(\theta^{\prime};Q)-R(\theta;Q)\bigl{\lVert}.$

### v3

@@ 5 Related Works
Continual Learning.
Early studies have documented catastrophic forgetting in both connectionist and backpropagation-based models, highlighting the fundamental stability-plasticity trade-off in continual learning (CL) (McCloskey and Cohen, 1989; Ratcliff, 1990). More recently, CL has been extended to LLMs, with methods such as rehearsal-based approaches (Robins, 1995; Lopez-Paz and Ranzato, 2017), parameter isolation techniques (Serra et al., 2018; Jung et al., 2020), and task arithmetic (Ilharco et al., 2022), primarily targeting the retention of task-specific knowledge. In contrast, SEAT addresses a different challenge: preserving abstention behavior on unseen queries during knowledge adaptation, a setting that does not map cleanly onto a standard task distribution.
Representation Learning.
Recent interpretability studies have shown that high-level cognitive phenomena in LLMs are encoded as linear directions in the model’s activation space (Park et al., 2023; Turner et al., 2023), and can be steered to encourage or suppress specific behaviors (Tian et al., 2025; Chen et al., 2025; Casademunt et al., 2025; Zou et al., 2023). SEAT takes a related but distinct perspective: rather than steering representations to induce a new behavior, it aims to limit unnecessary drift during fine-tuning so that existing abstention behavior is better preserved on unseen queries.
@@ 6 Conclusion
We study a practical failure mode of knowledge adaptation: fine-tuning can compromise an LLM’s ability to abstain on unsupported queries. Rather than presenting a formal guarantee, we develop an empirically grounded view in which loss of abstention behavior is associated with excessive representational drift and spillover to unseen neighboring entities. Guided by this perspective, we propose SEAT, a simple yet robust training recipe combining sparse updates and entity-perturbed regularization. Across the evaluated settings, SEAT consistently improves preservation of abstention behavior while maintaining strong performance on target knowledge updates. We hope this framing helps establish preservation of epistemic abstention as a useful objective for safer model adaptation.
@@ Appendix A Proofs
We show below the connection between sparse training, a core component of SEAT, and the constraint it imposes on the displacement of global residual stream activations. First, we present key properties of the residual map in Propositions 1 and 2, then we show sparse training constrains gradient norm which, in turn, constrains global residual stream activation movement.
Important to note that the discussion below highlights the role of sparse training in promoting global stability of residual stream activations during fine-tuning, rather than targeting epistemic abstention specifically. However, epistemic abstention is particularly sensitive to activation displacement, making such global stability critical for its preservation. Complemented by EP, which enhances local sharpness, both components are essential to the effectiveness of SEAT, as was discussed in the ablation study (§ 4.3).
Proposition 1.
@@
A decoder-only transformer model is a finite composition of primitives. Using Llama3 (Dubey et al., 2024) as a proxy, we list its modules, the formula implemented and its smoothness class below.
@@
Now. let $\mathcal{U}\subseteq\{1,\dots,P\}$ be the trainable coordinates and $\mathcal{F}=\mathcal{U}^{\mathrm{c}}$ be the frozen ones. Define sparse fine-tuning as $\theta^{\prime}\;=\;\theta\;-\;\eta\,M\,\nabla_{\theta}L(\theta)$, where $M$ is the mask matrix.
@@
Lemma 3 (Sparse fine-tuning constrains gradient-norm).
@@
Lemma 4 (Gradient-norm $\Rightarrow$ residual stream activation displacement).
For every layer $\ell$ and training step,
@@
Let $\gamma(t)=\theta+t(\theta^{\prime}-\theta)$ for $t\in[0,1]$. By Proposition 1, the residual map $\vartheta\mapsto R_{\ell}(\vartheta;x)$ is continuously differentiable on an open neighborhood of $K$, and by Proposition 2, $\gamma([0,1])\subset K$ and $K$ is compact. By the fundamental theorem of calculus for curves in $\mathbb{R}^{m}$
@@
Hence the supremum is $\leq L_{\ell}$. Finally $\|\theta^{\prime}-\theta\|=\eta\|\nabla_{\theta}\mathcal{L}(\theta)\|$, yielding the deterministic bound. ∎
By Lemma 3, we see that the masked gradient update used in sparse fine-tuning is non-expansive relative to the dense gradient update (i.e., $\|M\nabla_{\theta}\mathcal{L}(\theta)\|\leq\|\nabla_{\theta}\mathcal{L}(\theta)\|.$). By Lemma 4, residual stream activation displacement is controlled by the norm of the parameter update. Therefore, under the same learning rate, sparse fine-tuning yields a no-looser upper bound on layer-wise activation displacement than dense fine-tuning. Concretely, for the sparse update $\theta^{\prime}=\theta-\eta M\nabla_{\theta}\mathcal{L}(\theta),$, we have
$\|R_{\ell}(\theta^{\prime};x)-R_{\ell}(\theta;x)\|\leq\eta L_{\ell}(K)\|M\nabla_{\theta}\mathcal{L}(\theta)\|\leq\eta L_{\ell}(K)\|\nabla_{\theta}\mathcal{L}(\theta)\|.$
Hence sparsity acts as an explicit mechanism for tightening the worst-case bound on representation drift during fine-tuning.